OpenStudy (anonymous):

7xy + 4 for x= -5 and y = -2

OpenStudy (anonymous):

7(-5)=-35 -35(-2)=70, 70+4=74

jimthompson5910 (jim_thompson5910):

7xy + 4 7(-5)(-2) + 4 -35(-2) + 4 70 + 4 74 So 7xy + 4 when x= -5 and y = -2

OpenStudy (anonymous):

yes i got it right thanks

jimthompson5910 (jim_thompson5910):

congrats

OpenStudy (anonymous):

thnx